Differentiate the following with regard to x

y=ln( square toot X multiple e^x^2)


1
Expert's answer
2021-07-08T12:02:35-0400

"y=ln( \sqrt {x} \times e^{x^2})"


Let "u=\sqrt {x} \times e^{x^2}\implies y=ln(u)\implies {dy\over du}={1\over u}"


Let "f=\sqrt x=x^{1\over 2} \implies f'={1\over 2}x^{{1\over 2}-1}={1\over 2}x^{-{1\over 2}}"


Let "g=e^{x^2} \implies g'=2xe^{x^2}"


We then use product rule of differentiation to differentiate "u=\sqrt {x} \times e^{x^2}" with respect to "x"


"{du \over dx}=f'g+fg'"


"={1\over 2}x^{-{1\over 2}}e^{x^2}+\sqrt x(2xe^{x^2})"



We then use chain rule of differentiation to differentiate "y=ln( \sqrt {x} \times e^{x^2})" with respect to "x"


"{dy\over dx}={dy\over du}\times{du\over dx}"


"={1\over u}\times({1\over 2}x^{-{1\over 2}}e^{x^2}+\sqrt x(2xe^{x^2}))"


"={1\over \sqrt {x} \times e^{x^2}}\times({1\over 2}x^{-{1\over 2}}e^{x^2}+\sqrt x(2xe^{x^2}))"


"={1\over \sqrt {x} \times e^{x^2}}\times({1\over 2(\sqrt x)}e^{x^2}+\sqrt x(2xe^{x^2}))"

"={e^{x^2}\over 2xe^{x^2}}+{\sqrt x(2xe^{x^2}\over \sqrt x(e^{x^2})}"

"={1\over 2x}+2x"


"\therefore {dy\over dx}={1+4x^2\over 2x}"